OpenStudy (anonymous):

Verify the Identity: cot(x-pi/2)=-tanx Need help solving, will medal immediately. (:

OpenStudy (jdoe0001):

hint: \(\bf cot\left( x-\frac{\pi }{2} \right)\implies \cfrac{cos\left( x-\frac{\pi }{2} \right)}{sin\left( x-\frac{\pi }{2} \right)} \\ \quad \\ \cfrac{cos(x)cos\left(\frac{\pi }{2} \right)+sin(x)sin\left(\frac{\pi }{2} \right)}{sin(x)cos\left(\frac{\pi }{2} \right)-cos(x)sin\left(\frac{\pi }{2} \right)}\)

OpenStudy (anonymous):

@jdoe0001 I'm still confused as to how to verify the identity

OpenStudy (jdoe0001):

well... what's the \(cos\left(\frac{\pi }{2} \right)?\) what about the \(sin\left(\frac{\pi }{2} \right) ?\)

OpenStudy (anonymous):

The cos is 0 and the sin is 1, correct? @jdoe0001

OpenStudy (jdoe0001):

yes so... .one sec

OpenStudy (jdoe0001):

\(\bf cot\left( x-\frac{\pi }{2} \right)\implies \cfrac{cos\left( x-\frac{\pi }{2} \right)}{sin\left( x-\frac{\pi }{2} \right)} \\ \quad \\ \cfrac{cos(x)cos\left(\frac{\pi }{2} \right)+sin(x)sin\left(\frac{\pi }{2} \right)}{sin(x)cos\left(\frac{\pi }{2} \right)-cos(x)sin\left(\frac{\pi }{2} \right)}\implies \cfrac{cos(x)\cdot 0+sin(x)\cdot 1}{sin(x)\cdot 0-cos(x)\cdot 1}\implies ?\)

OpenStudy (jdoe0001):

what are you left with?

OpenStudy (anonymous):

We are left with sin(x)/-cos(x) which equals -tanx. Oh my gosh I get it thank you so much!

OpenStudy (jdoe0001):

yw